OpenStudy (unklerhaukus):

Solve the second order DE \[y^{\prime\prime}+4y=\cos2x\]

OpenStudy (unklerhaukus):

\[y_c^{\prime\prime}+4y_c=0\]\[m^2+4=0\]\[m=\pm2i\] \[y_c=A\cos2x+B\sin {2x}\]

OpenStudy (unklerhaukus):

\[y_p^{\prime\prime}+4y_p=\cos2x\] \[y_p=C\cos2x+D\sin2x\]\[y_p^\prime=-2C\sin2x+2D\cos2x\]\[y_p^{\prime\prime}=-4C\cos2x-4D\sin2x\]

OpenStudy (unklerhaukus):

or was i ment to multiply by x for some reason?

OpenStudy (fwizbang):

The particular solution your looking for is yp= Ax sin(2x) because you are driving the system at its resonant frequency(cos (2x) is a solution of the homogeneous equation.)

OpenStudy (unklerhaukus):

can you explain further

OpenStudy (shayaan_mustafa):

Hi Everyone! How are you all?

OpenStudy (unklerhaukus):

i was just following a table

OpenStudy (eyust707):

ohhh okay

OpenStudy (unklerhaukus):

hi there

OpenStudy (unklerhaukus):

do you know of a better table?

OpenStudy (shayaan_mustafa):

@UnkleRhaukus I want to say you that you must try again by your own self. When we try to solve any differential equation then complementary solution is much easy to find without any problem just like algebra. But for particular solution there are several methods to find this part. I recommend you to use "Variation of Parameters", the most powerful method to find particular solution.

OpenStudy (shayaan_mustafa):

Use Variation of Parameters. If you get command on this method then you will never get help for others.

OpenStudy (eyust707):

yeaa UNK you were right since the particular solution is already in the form of the of the complimentary you need to multiply by x

OpenStudy (eyust707):

\[y _{p}= x(Ccos(2x)+Dsin(2x))\]

OpenStudy (eyust707):

since Acos(2x) was already in the compliment

OpenStudy (eyust707):

Then you distribute the x. Take derivatives like you did before. There will be a few product rules. Plug the derivatives back in to the diff. You should get C =0 D = 1/4

OpenStudy (fwizbang):

Here's a quick and dirty derivation. Start with the problem \[y \prime \prime + 4 y = \cos (\Omega x)\] a particular solution(remember that there is an infinite set of particular solutions ,not just one) is then \[y _{p}= (\cos \Omega x -\cos 2x) / (4-\Omega^2)\] If \[\Omega = 2\] , this is indeterminate (0/0) so apply L'hopital's rule as \[ \Omega -> 2\] to get \[y _{p}= x \sin(2x)/4\]

OpenStudy (eyust707):

woah that method looks cool... In my diffs class we never learned that method.

OpenStudy (unklerhaukus):

\[y_p=x(C\cos2x+D\sin2x)\] \[y_p^\prime=x(-2C\sin2x+2D\cos2x)+C\cos2x+D\sin2x\] \[y_p^{\prime\prime}=(-2C\sin2x+2D\cos2x)+x(-4Cx\cos2x-4Dx\sin2x)\]\[\qquad\qquad-2C\sin2x+2D\cos 2x\]

OpenStudy (eyust707):

looks good...

OpenStudy (unklerhaukus):

\[y_p^{\prime\prime}=4(-C\sin2x+D\cos2x)+4x(-Cx\cos2x-Dx\sin2x)\]

OpenStudy (unklerhaukus):

\[\small(-4C\sin2x+4D\cos2x)+x(-4Cx\cos2x-4Dx\sin2x)+4x(C\cos2x+D\sin2x)=\cos2x\]

OpenStudy (unklerhaukus):

oh good, i might be ok from here, ill let you know if i get stuck again,

OpenStudy (eyust707):

yea these sometimes get hairy, haha gotta love diffs

OpenStudy (unklerhaukus):

\[(-4C-4Dx^2+4Dx)\sin2x+(4D-4Cx^2+4Cx)\cos2x=\cos2x\] \[(C+Dx^2-Dx)=0;\qquad (D-Cx^2+Cx)=1\] \[C+Dx(x+1)=0;\qquad D-Cx(x+1)=1\] \[D=1+Cx(x+1);\qquad C+\left(1+Cx\left(x+1\right)\right)x\left(x+1\right)=0\]

OpenStudy (unklerhaukus):

how am ment to solve for these constants?

OpenStudy (eyust707):

Usually a bunch of the terms cancel

OpenStudy (eyust707):

\[(-4C\sin2x+4D\cos2x)+x(-4Cx\cos2x-4Dx\sin2x)+4x(C\cos2x+D\sin2x)=\cos2x\]

OpenStudy (unklerhaukus):

\[\small(-4C\sin2x+4D\cos2x)+x(-4Cx\cos2x-4Dx\sin2x)+4x(C\cos2x+D\sin2x)=\cos2x\]

OpenStudy (unklerhaukus):

i dunno where to cancel

OpenStudy (eyust707):

there's probably a little error let me go back through real quick

OpenStudy (fwizbang):

You were good righht up to the check mark. There shouldn't be any terms with an x^2, just solve the equation you had at that point. (the coeffs of sin and cos need to vanish separately.)

OpenStudy (unklerhaukus):

\[\small(-4C\sin2x+4D\cos2x)+x(-4Cx\cos2x-4Dx\sin2x)+4x(C\cos2x+D\sin2x)=\cos2x\] silly me , this step is easy \[(-4C\sin2x+4D\cos2x)=\cos2x\]

OpenStudy (fwizbang):

It's always the little things that mess us up....

OpenStudy (eyust707):

\[y'' _{p} = -4Cxcos(2x) +4Dcos(2x) -4Dxsin(2x)-4Csin(2x)\]

OpenStudy (eyust707):

=)

OpenStudy (unklerhaukus):

\[-4C=0;\qquad 4D=1\]\[C=0;\qquad D=\frac 14\]\[y_p=\frac x4\cos2x\]

OpenStudy (fwizbang):

not quite.....

OpenStudy (unklerhaukus):

ah im out of of phase for some reason

OpenStudy (fwizbang):

go pack tp yp, what does D multiply?

OpenStudy (unklerhaukus):

oh haha

OpenStudy (unklerhaukus):

\[y_p=x(C\cos2x+D\sin2x)\] \[y_p=\frac x4\sin2x\]

OpenStudy (unklerhaukus):

\[y=y_c+y_p\] \[{y (x)= A\cos2x+B\sin2x}+\frac{1}4x\sin2x\quad \Large\color\red\checkmark\]

OpenStudy (unklerhaukus):

Thank you all ever so much ,

OpenStudy (eyust707):

YES!